Session: Sep 1999 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Thu Sep 30 02:24:59 METDST 1999) Success

Jodrell (Alastair Gunn at Thu Sep 30 10:32:37 METDST 1999) Success
No known problems.

Effelsberg (A. Kraus at Thu Sep 30 10:39:11 METDST 1999) Minor failures
We still have some problems with thin tapes on the VLBA4 recoreder (parity errors!). Reason unknown so far.

Noto (C.Trigilio at Thu Sep 30 18:42:52 METDST 1999) Success

Torun (K.Borkowski at Wed Oct 6 09:36:06 METDST 1999) Success
No known problems.

Onsala (A. Polatidis at Wed Oct 6 11:05:55 METDST 1999) Success

Calibration (Jean\-Francois Desmurs at Sun Nov 7 16:00:34 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: gah2.antab. Read the comments in the file to get further information (Rem : Wb missing)